OpenStudy (anonymous):

will medal and fan What property is shown by the equation? Column A 1. –24 • 0 = 0 2. 1/16x(16)=1 3. –18 • (–1) = 18 4. 5 ÷ (–1) = –5 Column B A. Reciprocal Property of Multiplication B. Identity Property of Division C. Negative One Property of Multiplication D. Zero Property of Multiplication E. Negative One Property of Division

OpenStudy (anonymous):

@haleyelizabeth

OpenStudy (anonymous):

@haleyelizabeth2017

OpenStudy (anonymous):

@richgall____

OpenStudy (anonymous):

@JosiahG

OpenStudy (anonymous):

Okay, so let's solve the first one together. What do you think the first one is based on its context?

OpenStudy (anonymous):

a

OpenStudy (anonymous):

What property states that any number multiplied by 0 is always 0?

OpenStudy (anonymous):

d

OpenStudy (anonymous):

So which number does that go along with?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Correct. Now number two says that 1/16 x 16/1 = 1. Which letter would that be?

OpenStudy (anonymous):

1d 2? 3? 4?

OpenStudy (anonymous):

c

OpenStudy (anonymous):

I would have to disagree with ya' on that one. Which property states that the opposite of a number multiplied by itself is 1?

OpenStudy (anonymous):

e

OpenStudy (anonymous):

We're not really dividing here, and it doesn't have a -1 in it. Which other one could it be?

OpenStudy (anonymous):

b

OpenStudy (anonymous):

Okay, it would be letter A and here's why: Reciprocal means that the opposite of a number multiplied by itself if 1. I'll show you the example: \[\frac{ 1 }{ 16 } * \frac{ 16 }{ 1 } = \frac{ 16 }{ 16 } = 1\]

OpenStudy (anonymous):

ok 1d 2a

OpenStudy (anonymous):

3? 4?

OpenStudy (anonymous):

Before I go further, are you understanding this or just looking for the answers?

OpenStudy (anonymous):

im understanding it i alreday had the frist one correct

OpenStudy (anonymous):

Okay, so number 3 would be the one that multiplied by -1 is either a double negative or the opposite of the number. Which property is that one?

OpenStudy (anonymous):

a

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

You're already stated that a is an answer. Which property states something about "-1"? Key word: -1

OpenStudy (anonymous):

You've*

OpenStudy (anonymous):

you can use it more then one time

OpenStudy (anonymous):

Even if you can, that's not the correct answer. Which property states about multiplying with negative one?

OpenStudy (anonymous):

oh d

OpenStudy (anonymous):

That states that any number multiplied by 0 = 0. 18 * (-1) = -18 Which property states about multiplying negative numbers?

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

e

OpenStudy (anonymous):

You're almost there, but that's division. We're doing multiplication.

OpenStudy (anonymous):

dang it

OpenStudy (anonymous):

Which one states about multiplying negative numbers?

OpenStudy (anonymous):

C!!!!!!

OpenStudy (anonymous):

Correct! Now I've noticed that number 4 is left out. What does number 4 state?

OpenStudy (anonymous):

5 ÷ (–1) = –5 e